Find the 43rd term of -124, -122, -120...
Answer by amarjeeth123(569)

The given series is in arithmetic progression.
First term a=-124
Common difference d=2
Nth term an=a+(n-1)d
a43=-124+(43-1)2=84-124=-40
Answer=-40